# Conformal Prediction Assessment: A Framework for Conditional Coverage Evaluation and Selection

### 28 Mar 2026 v1

@@ 21 $\hat{C}_{\text{new}}\leftarrow\mathcal{A}_{\text{final}}(X_{\text{new}})$
Appendix C Calibration of the Reliability Estimator
@@
C.1 The Role and Scope of Calibration in CPA
@@
C.2 Post-Hoc Calibration Procedure
@@
C.3 Diagnostic Metrics: Reliability Diagrams and ECE
@@
Appendix D Details of Data Generating Processes
@@
D.1 Setting A: Linear, Homoscedastic
@@
D.2 Setting B: Nonlinear, Heavy-Tailed
@@ D.3 Setting C: Heteroscedastic
D.4 Setting D: Heteroscedastic with Complex Covariates
@@
Appendix E Implementation Details of Benchmark Methods
@@
E.1 Implementation Details of Classical Statistical Baselines
@@
E.1.1 Ordinary Least Squares (OLS) for Standard Regimes
@@
E.1.2 Spline-Based Generalized Additive Models (GAM) for Nonlinearity
@@
E.1.3 Two-Stage Weighted Least Squares (WLS) for Heteroscedasticity
@@
E.2 Adaptive Residual Bootstrap
@@
E.3 Quantile Regression Forests (QRF)
@@
E.4 Split Conformal Prediction with Adaptive Selection
@@
E.5 Locally Adaptive (Studentized) Split Conformal Prediction
@@
E.6 Conformalized Quantile Regression (CQR)
@@
E.7 Cross-Validation Plus (CV+)
@@
E.8 Localized Conformal Prediction (LCP)
@@
E.9 Randomized Local Conformal Prediction (RLCP)
@@
Appendix F Definitions of Ranking Metrics
@@
F.1 Distance-Weighted Kendall’s Tau ($\tau_{w}$)
@@
F.2 Normalized Discounted Cumulative Gain (NDCG)
@@ F.3 Hit Rate (Hit@k)
Appendix G Details of Reliability Estimator Configurations
G.1 The Automated Model Selection Procedure (Baseline)
@@ G.2 Perturbation Configurations
Appendix H Bias Analysis and Mechanism of Failure
@@
Appendix I Sensitivity Analysis of Data Allocation
@@
I.1 The Bias-Variance Trade-off
@@ I.2 Empirical Results
Appendix J Empirical Investigation of Sample Complexity
@@
  • CPA Implementation: Following our heuristic recommendations, we employ a balanced split ($\rho=0.5$). The reliability estimator $\hat{\eta}(x)$ is trained via the AutoML pipeline described in Appendix G to ensure maximally flexible modeling of the coverage boundary.
@@
Appendix K Additional Details for Real-Data Experiments
K.1 Worst Slab Coverage
@@
K.2 Detailed Numerical Results for Real Data Experiments
@@
Method Selected Bootstrap CV+ CP-Residual CP-Studentized CQR LCP RLCP Dataset Metric Bike Length 161.847 230.635 217.013 234.722 182.445 156.291 128.791 138.572 Coverage 0.898 0.906 0.905 0.896 0.897 0.898 0.898 0.906 WSC-B 0.782 0.747 0.753 0.735 0.784 0.780 0.721 0.758 Computer Length 7.234 7.892 7.711 7.925 7.234 7.877 7.314 7.267 Coverage 0.895 0.900 0.898 0.895 0.895 0.900 0.897 0.910 WSC-B 0.769 0.704 0.683 0.669 0.769 0.753 0.718 0.774 Debutanizer Length 0.254 0.253 0.244 0.274 0.235 0.303 0.263 0.262 Coverage 0.911 0.910 0.909 0.903 0.905 0.906 0.900 0.916 WSC-B 0.680 0.635 0.653 0.611 0.670 0.637 0.625 0.668 Kin8nm Length 0.481 0.480 0.476 0.504 0.473 0.566 0.476 0.500 Coverage 0.899 0.904 0.905 0.904 0.898 0.901 0.898 0.901 WSC-B 0.765 0.742 0.744 0.740 0.761 0.772 0.720 0.739 Meps_21 Length 23.276 34.378 34.431 34.525 23.276 28.222 29.574 35.332 Coverage 0.902 0.902 0.904 0.902 0.902 0.955 0.904 0.899 WSC-B 0.804 0.478 0.457 0.442 0.804 0.883 0.518 0.457 Miami Length 0.446 0.484 0.483 0.498 0.446 0.523 0.469 0.500 Coverage 0.900 0.908 0.909 0.899 0.900 0.901 0.901 0.910 WSC-B 0.804 0.696 0.692 0.655 0.804 0.782 0.728 0.753 Parkinsons Length 16.992 17.877 17.866 19.300 16.946 17.379 11.335 11.249 Coverage 0.896 0.898 0.900 0.896 0.898 0.902 0.899 0.905 WSC-B 0.740 0.714 0.715 0.712 0.743 0.738 0.725 0.734 Qsar Length 2.665 2.688 2.640 2.820 2.672 3.467 2.830 2.861 Coverage 0.905 0.909 0.910 0.898 0.905 0.898 0.897 0.897 WSC-B 0.761 0.763 0.757 0.734 0.762 0.753 0.732 0.736 Temperature Length 2.428 2.453 2.431 2.604 2.443 3.010 2.341 2.716 Coverage 0.904 0.902 0.906 0.901 0.902 0.908 0.903 0.904 WSC-B 0.776 0.740 0.738 0.742 0.778 0.767 0.758 0.759
@@
K.3 Multivariate Reliability Landscapes for Bike Sharing Demand

### 23 Apr 2026 v2

@@ 21 $\hat{C}_{\text{new}}\leftarrow\mathcal{A}_{\text{final}}(X_{\text{new}})$
Appendix D Calibration of the Reliability Estimator
@@
D.1 The Role and Scope of Calibration in CPA
@@
D.2 Post-Hoc Calibration Procedure
@@
D.3 Diagnostic Metrics: Reliability Diagrams and ECE
@@
Appendix E Details of Data Generating Processes
@@
E.1 Setting A: Linear, Homoscedastic
@@
E.2 Setting B: Nonlinear, Heavy-Tailed
@@ E.3 Setting C: Heteroscedastic
E.4 Setting D: Heteroscedastic with Complex Covariates
@@
Appendix F Implementation Details of Benchmark Methods
@@
F.1 Implementation Details of Classical Statistical Baselines
@@
F.1.1 Ordinary Least Squares (OLS) for Standard Regimes
@@
F.1.2 Spline-Based Generalized Additive Models (GAM) for Nonlinearity
@@
F.1.3 Two-Stage Weighted Least Squares (WLS) for Heteroscedasticity
@@
F.2 Adaptive Residual Bootstrap
@@
F.3 Quantile Regression Forests (QRF)
@@
F.4 Split Conformal Prediction with Adaptive Selection
@@
F.5 Locally Adaptive (Studentized) Split Conformal Prediction
@@
F.6 Conformalized Quantile Regression (CQR)
@@
F.7 Cross-Validation Plus (CV+)
@@
F.8 Localized Conformal Prediction (LCP)
@@
F.9 Randomized Local Conformal Prediction (RLCP)
@@
Appendix G Definitions of Ranking Metrics
@@
G.1 Distance-Weighted Kendall’s Tau ($\tau_{w}$)
@@
G.2 Normalized Discounted Cumulative Gain (NDCG)
@@ G.3 Hit Rate (Hit@k)
Appendix H Details of Reliability Estimator Configurations
H.1 The Automated Model Selection Procedure (Baseline)
@@ H.2 Perturbation Configurations
Appendix I Bias Analysis and Mechanism of Failure
@@
Appendix J Sensitivity Analysis of Data Allocation
@@
J.1 The Bias-Variance Trade-off
@@
Appendix K Empirical Investigation of Sample Complexity
@@
  • CPA Implementation: Following our heuristic recommendations, we employ a balanced split ($\rho=0.5$). The reliability estimator $\hat{\eta}(x)$ is trained via the AutoML pipeline described in Appendix H to ensure maximally flexible modeling of the coverage boundary.
@@
Appendix L Additional Details for Real-Data Experiments
L.1 Worst Slab Coverage
@@
L.2 Detailed Numerical Results for Real Data Experiments
@@
Method Selected Bootstrap CV+ CP-Residual CP-Studentized CQR LCP RLCP Dataset Metric Bike Length 161.847 230.635 217.013 234.722 182.445 156.291 128.791 138.572 Coverage 0.898 0.906 0.905 0.896 0.897 0.898 0.898 0.906 WSC-B 0.782 0.747 0.753 0.735 0.784 0.780 0.721 0.758 Computer Length 7.234 7.892 7.711 7.925 7.234 7.877 7.314 7.267 Coverage 0.895 0.900 0.898 0.895 0.895 0.900 0.897 0.910 WSC-B 0.769 0.704 0.683 0.669 0.769 0.753 0.718 0.774 Debutanizer Length 0.254 0.253 0.244 0.274 0.235 0.303 0.263 0.262 Coverage 0.911 0.910 0.909 0.903 0.905 0.906 0.900 0.916 WSC-B 0.680 0.635 0.653 0.611 0.670 0.637 0.625 0.668 Kin8nm Length 0.481 0.480 0.476 0.504 0.473 0.566 0.476 0.500 Coverage 0.899 0.904 0.905 0.904 0.898 0.901 0.898 0.901 WSC-B 0.765 0.742 0.744 0.740 0.761 0.772 0.720 0.739 Meps_21 Length 23.276 34.378 34.431 34.525 23.276 28.222 29.574 35.332 Coverage 0.902 0.902 0.904 0.902 0.902 0.955 0.904 0.899 WSC-B 0.804 0.478 0.457 0.442 0.804 0.883 0.518 0.457 Miami_2016 Length 0.446 0.484 0.483 0.498 0.446 0.523 0.469 0.500 Coverage 0.900 0.908 0.909 0.899 0.900 0.901 0.901 0.910 WSC-B 0.804 0.696 0.692 0.655 0.804 0.782 0.728 0.753 Parkinsons Length 16.992 17.877 17.866 19.300 16.946 17.379 11.335 11.249 Coverage 0.896 0.898 0.900 0.896 0.898 0.902 0.899 0.905 WSC-B 0.740 0.714 0.715 0.712 0.743 0.738 0.725 0.734 Qsar Length 2.665 2.688 2.640 2.820 2.672 3.467 2.830 2.861 Coverage 0.905 0.909 0.910 0.898 0.905 0.898 0.897 0.897 WSC-B 0.761 0.763 0.757 0.734 0.762 0.753 0.732 0.736 Temperature Length 2.428 2.453 2.431 2.604 2.443 3.010 2.341 2.716 Coverage 0.904 0.902 0.906 0.901 0.902 0.908 0.903 0.904 WSC-B 0.776 0.740 0.738 0.742 0.778 0.767 0.758 0.759
@@
L.3 Multivariate Reliability Landscapes for Bike Sharing Demand